OpenStudy (anonymous):

what values for theta(0 equal/less than theta equal/less than 2 pi) satisfy the equation? 2 sin theta cos theta + sqrt 3 cos theta=0

OpenStudy (anonymous):

what values for $$\theta$$ if \[0 \le \theta \le 2\pi\] satisfy the equation $$2sin(\theta)cos(\theta) + \sqrt{3} cos(\theta) = 0$$

OpenStudy (anonymous):

How do you do that?

OpenStudy (anonymous):

you can click equation button for easy editing or use latex manually http://en.wikibooks.org/wiki/LaTeX/Mathematics

OpenStudy (anonymous):

These are my choices, and I can't seem to figure out how to use that website.

OpenStudy (anonymous):

since you are given choices all you have to do is plug in the angles and see if it equals 0

OpenStudy (anonymous):

I use them in place of theta? How come there's four numbers, but only three thetas in the equation?

OpenStudy (anonymous):

yes replace all thetas in the equation with one number in the comma separated list, then replace all thetas int he equation with the next number in the comma separated list, etc

OpenStudy (anonymous):

because there are four values (or angles) for theta that will make that equation true, on the interval 0 <= theta <= 2pi

OpenStudy (anonymous):

if you plug one in and get a number not equal to 0, move on to the next list

OpenStudy (anonymous):

None of them equal 0

OpenStudy (anonymous):

yes several of them do, if you plug in pi/2 for theta it will equal 0, so unfortunately since that is in all 4 lists you cant eliminate any lists

OpenStudy (anonymous):

so move on to the next value in the first list

OpenStudy (anonymous):

yep, that equals 0, without the parenthesis its a little unclear but yeah that equals 0

OpenStudy (anonymous):

When I type that into google looking for an answer, it gives me this -0.86602540378

OpenStudy (anonymous):

like i said you really need parenthesis, try this 2sin(pi/2)cos(pi/2)+sqrt(3)cos(pi/2)

OpenStudy (anonymous):

It's the first set of values, right?

OpenStudy (anonymous):

that looks good to me